OpenStudy (anonymous):

Calculate the consumers' surplus at the indicated unit price (P with a bar over it) for the demand equation. q=20-0.05p^2; (P with bar over it) = 2

OpenStudy (turingtest):

nobody knows if they can help you if you don't post your actual question ;)

OpenStudy (anonymous):

good point haha my first question is Calculate the consumers' surplus at the indicated unit price (P with a bar over it) for the demand equation. q=20-0.05p^2; (P with bar over it) = 2

OpenStudy (anonymous):

I can send you some examples if you need them to get your bearings

OpenStudy (turingtest):

maybe that would help, this seems to require some context just so you know, it's better to post the question like that in the original post the problem here though is terminology... p is your "unit price" if so I would think you just plug in the number 2 for p and get your answer q, but you say this is a calculus problem?

OpenStudy (anonymous):

Yeah its considered calculus in my book the chapter is labeled "Further integration techniques and Applications of the integral"

OpenStudy (anonymous):

let me find you an example one moment

OpenStudy (turingtest):

so then you must be integrating with respect to p over some interval?

OpenStudy (anonymous):

Yes exactly! :D

OpenStudy (anonymous):

ok just a few seconds more waiting for picture of example to save

OpenStudy (turingtest):

there is a button that says "equation" below on most browsers that lets you type the equation in

OpenStudy (turingtest):

\[\int_a^b f(x)dx\]

OpenStudy (anonymous):

its backwards but the answer is 40.5

OpenStudy (anonymous):

the issues I am having with the current problem however though is q equals the equation instead of p which is what i'm used to doing.

OpenStudy (turingtest):

in this case the "unit price" went from 0 to 27 or what? where did the bounds come from?

OpenStudy (turingtest):

all this depends on what these variables represent; we need more context here...

OpenStudy (anonymous):

yes the unit price goes from 0 to 27 which is why we use the interval 0 to 27

OpenStudy (turingtest):

and what are you calculating in that paper?

OpenStudy (anonymous):

The consumers' surplus which is basically how much a customer is willing to pay for a given item

OpenStudy (turingtest):

and where does the 4 come from?

OpenStudy (turingtest):

you say p=4 in the paper? is that like p=2 in this question?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

you take the p bar value and substitute it in for the value of p

OpenStudy (turingtest):

ok, so this problem is the exact same as the other one, but with p=2, and you don't know the bounds? you are calculating the same thing you say; the consumer price... thing

OpenStudy (turingtest):

oh I see

OpenStudy (anonymous):

so if I were to do it my q would end up equalling 79.8 but thats just the first part

OpenStudy (anonymous):

after obtaining the q value we still have to integrate it

OpenStudy (turingtest):

but in your example, the integration you do results in q?

OpenStudy (anonymous):

so I know we would use the integral \[\int\limits_{0}^{27}\]

OpenStudy (anonymous):

but I am not sure what we would use as the equation to put inside that integral

OpenStudy (anonymous):

if you look closely at the example our original equation is p=10-2q^(1/3) after substituting in 4 for p and subtracting 10 from either side we are left with -6=-2q^1/3

OpenStudy (turingtest):

oh man I really wish that paper hadn't been backwards, I didn't realize there were q's in there the whole time :P

OpenStudy (anonymous):

which we then later use in our interval as \[\int\limits_{0}^{27}(6-2q^(1/3)-4dq\]

OpenStudy (anonymous):

hahaha Its all good that one is on me xD

OpenStudy (anonymous):

after integrating we write (6-2q^(4/3))-4q) still using the interval 0 to 27

OpenStudy (turingtest):

I don't really knw what you're doing, or where you get 27 from again, but if I had to guess I would think you'd just turn it around\[q=20-0.5p^2\implies p=\sqrt{40-2q}\]

OpenStudy (turingtest):

and integrate that

OpenStudy (turingtest):

I'm sorry, for one thing this is talking about I subject I don't know. math yes, business or econ or whatever this is, no.... another is I'd really like to at least have that pic not reversed.

OpenStudy (anonymous):

so the 27 comes from us taking -6=2q^(1/3) which we got from substituting in 4 for p in the original equation. We then have to isolate the q in -6=-2q^(1/3) we do that by dividing each side by -2

OpenStudy (anonymous):

give me one moment and I'll have a non-reversed version for you :)

OpenStudy (turingtest):

actually I have to go, but I suggest you repost your question along with this prior explanation (not reversed of course) and someone will be able to help you. Good luck

OpenStudy (anonymous):

Thank you have a good evening:)